OpenStudy (anonymous):

perform the operation and write it in standard form. Please help! (i/3-2i)+(2i/3+8i)

OpenStudy (primeralph):

Just add normally.

OpenStudy (anonymous):

i would write both in standard form first, then add \[\frac{i}{3-2i}=\frac{i}{3-2i}\times \frac{3+2i}{3+2i}=\frac{-2+3i}{9+4}=\frac{-2}{13}+\frac{3}{13}i\]

OpenStudy (primeralph):

It looks different in the prompt. Which one is it? i/(a+bi) or i/a+bi?

OpenStudy (anonymous):

you have to do the same thing with the other one then you can add up

OpenStudy (anonymous):

i was adding first which was why i got confused but thank you!!

OpenStudy (anonymous):

yw